OpenStudy (anonymous):

find the slope of the line: 2(1-y)=x

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

do you need help with that?

OpenStudy (abb0t):

distribute the 2 across the (1-y)

OpenStudy (anonymous):

i got y=x+2 am i wrong?

OpenStudy (anonymous):

almost,like abbb0t said distribute the 2 across by dividing both sides by two

OpenStudy (anonymous):

then you have 1-y=(1/2)x

OpenStudy (anonymous):

Can you solve for y from here?

OpenStudy (anonymous):

i'll try. i'll let you know what i get.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

i got y=-(1/2)x+1

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

so you know how to get the slope from here?

OpenStudy (anonymous):

yes. y=mx+b :)

OpenStudy (anonymous):

cool :)